An out-of-control 320 kg boat speeds at 15 m/s towards shore where several small children are playing. Wonder Woman steps in to stop the boat. If on average she can apply 1500 N of force, how long does it take her to stop the boat?
Answer #1

given mass m = 320 kg

speed v = 15 m/s

applied force f = 1500 N

from impulse momentum theorem

impulse = change in momentum

impulse = f*t

f*t = m*v

t = m*v/f

t = 320*15/1500

t = 3.2 sec